### v1

# Diffusion-VLA: Scaling Robot Foundation Models via Unified Diffusion and Autoregression

Abstract
In this paper, we present DiVLA, a novel framework that seamlessly combines the autoregression model with the diffusion model for learning visuomotor policy. Central to our approach is a next-token prediction objective, enabling the model to reason effectively over the user’s query in the context of current observations. Subsequently, a diffusion model is attached to generate robust action outputs. To enhance policy learning through self-reasoning, we introduce a novel reasoning injection module that integrates reasoning phrases directly into the policy learning process. The whole framework is simple and flexible, making it easy to deploy and upgrade.
We conduct extensive experiments using multiple real robots to validate the effectiveness of DiVLA. Our tests include a challenging factory sorting task, where DiVLA successfully categorizes objects, including those not seen during training. We observe that the reasoning module enhances interpretability, allowing observers to understand the model’s thought process and identify potential causes of policy failures. Additionally, we test DiVLA on a zero-shot bin-picking task, achieving 63.7% accuracy on 102 previously unseen objects. Our method demonstrates robustness to visual changes, such as distractors and new backgrounds, and easily adapts to new embodiments. Furthermore, DiVLA can follow novel instructions and retain conversational ability. Notably, DiVLA is data-efficient and fast at inference; our smallest DiVLA-2B runs 82Hz on a single A6000 GPU and can train from scratch on less than 50 demonstrations for a complex task. Finally, we scale the model from 2B to 72B parameters, showcasing improved generalization capabilities with increased model size.
1 Introduction
Autoregressive models are predominant in Large Language Models, operating by sequentially predicting discrete tokens, where each token generation depends on the preceding ones. This approach has also influenced robotics, leading to the development of Vision-Language-Action (VLA) models, which frame action prediction as a next-token prediction task. While these models, i.e., RT-2 [9] and OpenVLA [26], have achieved notable success, they face intrinsic limitations. First, discretizing continuous action data into fixed-size tokens disrupts the coherence and precision of actions. Second, NTP is inherently inefficient for action generation, especially in real-time robotic applications where performance is critical.
Meanwhile, building on the success of diffusion models in content generation [43, 36, 38, 15, 16], diffusion-based models for learning visuomotor policies [10] have gained significant popularity over the past two years. Numerous methods have demonstrated strong performance in manipulation tasks by modeling action sequence generation as a noise-denoising process. This approach better captures the multimodal nature of robotic actions and enables faster sequence generation compared to NTP-based VLA models. However, despite the advantages of diffusion models for policy learning, they lack the reasoning capabilities crucial for VLA models to solve complex tasks effectively, a component that evidently improves LLMs.
This motivates us to raise the question: can we bring together the advantages of both parties, specifically the reasoning power of autoregressive models and the robustness of high-frequency action generation offered by diffusion models?
In this work, we propose a unified model, named DiffusionVLA (DiVLA in short), that integrates autoregression with a diffusion model. The autoregressive component provides reasoning over the query, while the diffusion model controls the robot. Specifically, DiVLA builds upon a pre-trained Vision-Language Model (VLM), retaining its autoregressive capabilities for text-based reasoning. We extend this foundation by integrating a diffusion model that facilitates learning robotic actions through a noise-denoising process. This setup empowers DiVLA to achieve both language-driven reasoning and robust action generation in robotic contexts. However, simply combining these elements does not fully exploit the reasoning potential, as there is often an implicit gap between logical reasoning and actionable robot policies. To bridge this gap, we propose a reasoning injection module, which reuses reasoning outputs and embeds them directly into the policy head, thus enriching the policy learning process with explicit reasoning signals. This innovation allows us to directly incorporate reasoning into action generation, enhancing the model’s dexterity, robustness, and generalization across various scenarios. Our experiments confirm that DiVLA delivers the following advantages:
  • Fast Inference Speed: DiVLA-2B achieves an inference rate of 82Hz, and DiVLA-7B operates at 42Hz on a single A6000 GPU, ensuring real-time responsiveness in high-demand environments.
  • Enhanced Visual Generalization: DiVLA is unaffected by visual distractions or novel backgrounds, demonstrating robustness in visually dynamic settings.
  • Generalizable Reasoning Capability: DiVLA can recognize and categorize previously unseen objects accurately, showcasing its ability to generalize reasoning across novel inputs.
  • Adaptability to Novel Instructions and Conversational Capability: Our approach can interpret and execute complex, novel instructions while maintaining conversational fluency, offering a versatile response range in interactive scenarios.
  • Generalization to Other Embodiment: DiVLA readily fine-tunes for deployment on real bimanual robots, achieving high performance with minimal adjustment, proving its adaptability across diverse robotic embodiments.
  • Scalability: We provide a scalable family of models — DiVLA-2B, 7B, and 72B — demonstrating that generalization and performance improve with model size, adhering to established scaling laws.
We believe that DiVLA offers a novel perspective on the learning and generalizing robot policy.
2 Related Works
Autoregression models. Predicting the next token has been regarded as a key approach toward general artificial intelligence due to its success in training language models [48, 49, 1, 4, 47]. RT-2 [9] pioneered the application of next-token prediction in robot learning which predicts actions by converting continuous actions into discrete tokens for learning robotic motion. Building on this, OpenVLA [26] introduced an open-source, improved, and smaller version of RT-2 [9] with a similar approach, while ECoT [64] developed a chain-of-thought method. However, research has shown that next-token prediction may not be the optimal approach for robot models, especially when adapting to various embodiments. In this work, we leverage the strength of next-token prediction specifically for reasoning tasks.
Diffusion models. Diffusion models have become dominant in the field of visual generation. The Diffusion Policy [10] extends the application of diffusion models to robot learning, demonstrating their effectiveness in handling multimodal action distributions. Subsequent work has advanced the Diffusion Policy [71, 55, 39, 42, 50, 51, 5, 6, 11, 30] by applying it to 3D settings [24, 68, 66, 62], scaling it up [74], increasing its efficiency [21, 56], and introducing architectural innovations. For instance, TinyVLA [57] integrates diffusion models with lightweight vision-language models, while $pi_{0}$ [7] leverages flow matching rather than diffusion for action generation. Our approach introduces reasoning—a crucial element in language models—into the diffusion-based vision-language-action (VLA) model.
Robot foundation models. Existing works [3, 8, 9, 14, 22, 20, 12, 75] leverage RL [63, 41, 19, 32, 45, 67, 65] and LLM [29, 58, 73, 2, 40, 18, 23] to decouple the multimodal understanding and embodied control. Another line of research leverages pre-trained Vision-Language Models (VLMs) and fine-tunes them directly on robotic data [9, 64, 26, 57]. Our work follows this approach, unifying autoregressive and diffusion models for both reasoning and manipulation tasks.
Unified auto-regressive model and image generation. Recent work has focused on unifying multimodal understanding with image generation. These efforts include using auto-regressive methods to generate images [28, 46, 54, 52, 33], diffusion models to produce text, or combining both approaches into unified models [13, 31, 70, 59], such as Show-O [61], Transfusion [72], and Vila-U [60]. Unlike these methods, our framework explores unifying next-token prediction with diffusion models, which enhances the robot model’s reasoning abilities. This reasoning capability, in turn, improves the model’s generalization to various tasks and environments.
3 Methodology
Our ultimate goal is to create a unified framework that combines autoregressive models, which excel at predicting language sequences for reasoning, with diffusion models, which are highly effective at generating robotic actions. Developing such an integrated model presents substantial challenges, with key issues centered on: (i) designing an architecture that seamlessly and efficiently integrates both autoregressive and diffusion mechanisms; and (ii) leveraging self-generated reasoning to enhance action generation without adding inference computation overhead. In this section, we introduce the overall framework of our method in Section 3.1 and explore the design choices that inform our model architecture in Section 3.2.
3.1 Architecture
Given any sequence of interleaved images, text, and video, we first encode the images into dense visual features using SigLIP [69]. These encodings are then transformed into a fixed number of $N$ visual embeddings through a Transformer. It’s worth noting that typical visual inputs in robot learning often include multiple camera views. To manage this, we applied the shared SigLIP visual backbone to each view and subsequently concatenated the resulting visual tokens.
For vision-language processing, we leveraged Qwen2-VL [53], a state-of-the-art vision-language model available in three sizes: 2B, 7B, and 72B parameters. We initialized the VLM backbone with the publicly released checkpoint. It is also possible to use any other pre-trained VLM as backbone, since we decouple the vision-language understanding with action generation, making the overall architecture flexible to fit for advanced new models.
Action decoder. We employ a latent diffusion model to decode visual and instruction embeddings into actions. Specifically, we generate a set of tokens using an LLM and feed these tokens into the diffusion model as conditioning inputs for action decoding. Our architecture follows the design of the standard Diffusion Policy [10], with randomly initialized model weights. This component also incorporates reasoning from the LLM, which we describe in detail below. An MLP layer is attached to the last layer at the bottom of the action decoder to predict the robot’s joint space. If multiple embodiments are evolved, instead of making a copy of a separate action decoder [34], we simply initialized a new MLP layer for training and evaluation. This step ensures that the knowledge from the pre-trained data is preserved and thus can quickly adapt to a new embodiment.
Reasoning injection module. The core of our approach lies in introducing explicit reasoning into Vision-Language-Action (VLA) models. Unlike most autoregressive VLAs, which require a recursive setup — converting reasoning outputs into inputs for subsequent model runs — our method proposes a more efficient and streamlined integration of reasoning. By embedding reasoning directly within the policy model, we avoid the computational and operational complexities of iterative input-output cycles, enabling faster and more seamless reasoning incorporation.
Our reasoning injection module operates by taking the final embedding from the tokenized output of the reasoning component and directly injecting it into the policy model through Feature-wise Linear Modulation (FiLM) [37]. This injection technique, inspired by methods in RT-1 [8] and YAY [44], enables us to modulate the policy network’s layers based on the reasoning signal. We refer to this process as “injection” because, in our design, the policy network focuses primarily on action-specific tokens, while the reasoning module functions as an auxiliary enhancement, providing contextual depth without dominating the primary decision-making flow. This approach ensures that reasoning is not only present but actively utilized during policy model training.
3.2 Model Design Choices
We illustrate the training strategy and other techniques that we used to improve the efficiency and effectiveness of our method. We also discuss the selection of pretraining data.
View-adaptive tokenization. As experiments in robot policy learning commonly use multiple camera views. As the number of views increases, the count of visual tokens also rises, leading to significant computational overhead due to the complexity of self-attention. To address this, we propose a view-adaptive tokenization method that reduces the number of visual tokens, optimizing computational efficiency.
Our approach is straightforward yet effective: we significantly reduce the number of tokens for wrist camera views to just 16, which is over 20 times fewer than those for extrinsic views. This reduction is motivated by the distinct role of the wrist camera, which primarily assists in precise localization for grasping tasks. High-resolution processing of wrist views, while detailed, is often inefficient and consumes resources that could be better allocated to processing more critical visual signals from external perspectives.
This optimization is particularly beneficial in dual-arm robotic systems, where two wrist cameras are active simultaneously. By reducing the computational load associated with wrist views, we free up resources to handle other essential tasks, enabling smoother, more efficient operation and potentially enhancing the overall responsiveness and precision of the robotic system.
Training objectives. Given a batch of input sequences, the overall training loss is formulated as a combination of the diffusion loss and the next-token prediction loss: $L=L_{diff}+\alpha L_{ntp}$, where $\alpha$ is the hyper-parameters weighting the loss term $L_{ntp}$. In our observations, the magnitude of $L_{ntp}$ consistently remains about ten times smaller than that of $L_{diff}$. To balance the contribution of each component to the overall loss, we typically set $\alpha=1$ in all experiments. This adjustment ensures that both terms are comparably weighted in the training process, allowing the model to learn effectively from both action and next-token prediction tasks.
Pretraining Data. We consider OXE [35] and Droid [25] dataset for pretraining. We use Droid data to pre-train DiVLA-2B and DiVLA-7B. For DiVLA-72B, we use OXE and Droid together for pre-train. The original Droid data contains only robotic actions, paired partially with observations and language instructions. These data contain only robotic actions, paired partially with observations and language instructions. To enhance our model’s ability to generalize with language, we leverage GPT-4o to automatically transform these data into a form that includes reasoning. Consequently, the network architecture remains consistent across both the pre-training and fine-tuning stages.
Figure 1: Environmental setup for the Franka Robot and experimental configuration for sorting/bin-picking. Left: For factor sorting tasks, (a) The target sorting box is divided into four distinct sectors, each designated for one of the following categories: stuffed toys, hex keys, knit gloves, and toy cars, (c) The seen objects in the train data, (d) mixing the seen and unseen object for evaluation, (e) cluttered scene for seen objects, (f) cluttered scene for mixing seen and unseen objects. Middle: We use a Franka robot arm equipped with two external Zed cameras and a Realsense 435i wrist camera. Right: The setup for zero-shot bin picking.
4 Experiments
In this section, we examine the effectiveness of DiVLA for embodied control. In Section 4.2, we compare DiVLA  against other state-of-the-art models within a standard multi-task setting, assessing its performance in both in-distribution and out-of-distribution scenarios. In Section 4.3, we evaluate DiVLA  in the challenging sorting task, showcasing its remarkable performance and illustrating how reasoning enables the model to analyze robot actions and sort items accurately. In Section 4.4, we showcase DiVLA’s impressive generalization abilities in a zero-shot bin-picking task involving over 102 unseen objects. In Section 4.5, we illustrate DiVLA’s adaptability to new embodiments, successfully completing complex tasks that require bimanual coordination. We provide experimental setup and implementation details in the Appendix.
4.1 Experimental Setup.
Implementation details and pretraiend data. The model is pre-trained on the Droid [25] dataset. We then finetune our model on evaluation tasks, similar to the setting as $\pi_{0}$ [7]. We use LoRA [17] to fine-tune the VLM models. We use 2e-5 as a fixed learning rate to pretrain the model.
Data for finetuning. We explore four experimental settings: sorting, bin picking, multi-task learning, and table bussing. The first three settings are conducted with the Franka robot, while the table bussing task utilizes the bimanual AgileX robot. Our dataset includes 500 trajectories for the sorting task and 580 trajectories for multi-task learning. The bin picking task is designed as a zero-shot task, so no training data was collected for it. For the table bussing task, we gathered 400 trajectories, where objects are randomly placed on the table, often overlapping with each other. During the fine-tuning stage, all data corresponding to the same embodiment are trained together. For instance, the sorting and multi-task data are combined for training purposes.
Figure 2: Experimental results for sorting. We compared our DiVLA with Diffusion Policy, Octo, TinyVLA, and OpenVLA. DiVLA achieves the highest average success rate, outperforming the runner-up OpenVLA by 20.9%.
Figure 3: Zero-shot bin picking on 102 unseen objects. Our method outperforms the state-of-the-art robot foundation models by a large margin.
Table 1: Experimental results for multi-task learning on real robot. We report the count of pre-trained trajectories. We also report the average success rate for evaluation on both in-distribution and out-of-distribution. Task 1: Select the appropriate object based on the user’s intent. Task 2: Flip the vertically placed pot. Task 3: Pick up the cube and place it into the [yellow/blue] box. Task 4: Place the cup onto the plate. Task 5: Place the cube into the box.
Pre-trained In-Distribution Visual Generalization
Model $\setminus$ Tasks Trajectory Task 1 Task 2 Task 3 Task 4 Task 5 Avg. Task 1 Task 2 Task 3 Task 4 Task 5 Avg.
Diffusion Policy [10] - 66.7 36.4 0 36.4 0 27.9 11.1 11.1 0 22.2 0 8.9
ScalingDP-1B [74] - 66.7 36.4 27.3 37.0 18.2 35.2 33.3 33.3 0 22.2 22.2 22.2
TinyVLA [57] - 72.7 45.5 36.4 45.5 27.3 45.5 44.4 44.4 11.1 22.2 22.2 28.9
Octo [34] 970K 57.6 27.3 9.1 0 27.3 24.3 44.4 11.1 11.1 0 22.2 17.8
OpenVLA-7B [26] 970K 69.7 18.2 18.2 36.4 54.5 39.4 55.6 11.1 0 33.3 33.3 26.7
DiVLA-2B 39K 100 100 63.6 63.6 90.9 83.6 44.4 66.7 44.4 66.7 66.7 57.8
DiVLA-7B 39K 100 100 72.7 81.8 100 90.9 66.7 88.9 55.6 77.8 66.7 71.0
Figure 4: Examples of visual variations, including randomly placed distractors, distracting lighting, and a colorful background on the right. DiVLA is robust to visual changes in different scenarios.
4.2 Real-World Multi-Task Learning
We begin with a standard setting in which the model is trained on multiple tasks and completes each task based on different user queries. We designed five tasks: object selection, flip the vertically placed pot, placing a cube into a designated box, placing a cup onto a plate, and placing a cube inside a box. Detailed descriptions of these tasks are provided in the Appendix. The experimental results can be found in Table 1.
We compare our method to the Diffusion Policy [10], TinyVLA [57], Octo [34], and OpenVLA [26]. Notice that both Octo and OpenVLA is pre-trained on OXE [35], which is 25 times larger than our pre-trained datasets.
Generalization to visual changes. We further evaluate our method in a multi-task setting with visual changes to assess its robustness and adaptability in diverse, dynamic environments. Specifically, we introduce three challenging scenarios designed to test the model’s ability to handle visual variability: 1) adding additional distractors in the surroundings to increase visual clutter and complexity, 2) altering the background to test resilience against shifts in scene context, and 3) implementing colorful lighting effects to introduce varied illumination and color hues. These scenarios are shown in Figure 4 to illustrate the impact of each change on the visual environment. The experimental results are demonstrated in Table 1.
Our evaluation of these scenarios reveals that while all methods experience a decline in performance due to these visual changes, our method consistently maintains the highest average success rate across five different tasks. This outcome highlights the model’s inherent robustness and adaptability, despite the absence of any specific data augmentation techniques during training.
4.3 End-to-End Sorting on Real Robot
We evaluate the capability of DiVLA in an industrial setting, where a robot is tasked with sorting items into designated sectors within a large box based on object category. Specifically, we categorize items into four classes: (1) toy cars, (2) knit gloves, (3) stuffed toys, and (4) hex keys. The language instruction provided is ”Sort all items into corresponding areas”. A total of 500 trajectories are collected as training data. The task is considered successful only if the robot successfully grasps the object and places it in the correct sector. The experimental setup is illustrated in Figure 1.
This task poses several challenges, requiring both precise object grasping and accurate category identification. We assess our method under two difficulty settings: easy and hard. In easy mode, fewer than 5 items are placed on the table, whereas in hard mode, 6 to 11 items are randomly arranged. Furthermore, both seen and unseen objects are mixed in these scenarios. In the cluttered scene, items may overlap or be randomly distributed across the table, increasing the complexity of the sorting task.
The experimental results are illustrated in Figure 2. DiVLA demonstrates robust performance with an average success rate of 66.2% across all experimental settings. While other methods show significant performance degradation as scene complexity increases (i.e., higher object count and clutter level), particularly evident in DP’s sharp decline to 9.2% success rate in highly cluttered mixed scenarios, DiVLA maintains a substantial 60% success rate. This sustained performance underscores our approach’s capability to effectively handle complex and dynamic real-world scenarios.
Diagnosing policy model by checking reasoning. Because our model generates output with natural language reasoning, we can gain insight into what the model is “thinking” by observing its reasoning phrases. For example, as shown in Figure 5, the model identifies a toy car and decides to pick it up. If we intervene by placing a hex key in the gripper instead, the reasoning phrase shifts from “grabbing the toy car” to “grabbing the hex,” allowing the model to adapt and sort the item accurately. This dynamic reasoning makes the model’s decision-making process more transparent and interpretable. The reasoning injection module also benefits from the reasoning self-correction such that the robot action is more robust.
Identifing and sorting unseen objects. We observe that our model can identify unseen objects by name. When we place four previously unseen objects on the plate, a stuffed toy cat, a pair of green gloves, a dark toy car, and a screwdriver. The model successfully identifies the first two objects as “brown toy cat” and “green gloves”, and then correctly sorts them. For the remaining two objects on the right, which it cannot recognize, the model labels them as “the black object” and “the object”. Although it fails to sort these unrecognized items, our model still demonstrates generalization in identifying objects and effectively translates these identifications into appropriate actions.
Figure 5: DiVLA adjusts its output reasoning based on changes to the target object.
Figure 6: Some of the unseen objects used for evaluation in the zero-shot bin-picking tasks.
Figure 7: Examples of various unseen objects in zero-shot bin-picking tasks. The unseen objects vary significantly in size, demonstrating that DiVLA  can successfully pick objects of different sizes without prior training on them.
Table 2: Experimental Results for Following Novel Instruction. We report the success rate, calculated as the number of successful trials divided by the total trials, for the novel instruction-following abilities of OpenVLA and DiVLA-2B. Our method successfully picks up items based on given instructions and can follow sequential instructions to retrieve items in the correct order.
Following Novel Instruction on Four Tasks
Model $\setminus$ Tasks Watermelon Watermelon $\rightarrow$ Blue Paper Trash $\rightarrow$ Lemonade
OpenVLA [26] 2/3 0/3
DiVLA-2B 2/3 2/3
[…]
OpenVLA [26] 0/3 0/3
DiVLA-2B 1/3 1/3
4.4 Zero-Shot Bin Picking of Unseen Objects
This section evaluates instance generalization for DiVLA, focusing specifically on the bin-picking task—a standard benchmark for assessing robot model performance. In our evaluation, we use 102 unique objects, none of which were included in the training data. Some of these objects are shown in Figure 6. We modify the task instruction to “move any object on the right panel to the left basket.” Figure 1 (right) illustrates the experimental setup. We show some of the objects used in this test in Figure 6. The challenge in this evaluation lies in the significant variability across objects, which includes not only differences in dimensions but also distinct color patterns, textures, and degrees of deformability. Such variability aims to emulate real-world scenarios where robots must adapt to diverse and unpredictable items. Figure 7 provides examples of five differently sized objects from this experiment.
The experimental results are depicted in Figure 3. DiVLA achieves a success rate of 63.7%. In comparison, the success rates of the Diffusion Policy, Octo, TinyVLA, and OpenVLA are 8.9%, 19.6%, 23.5%, and 28.4%, respectively. These results indicate that DiVLA is resilient across a broad range of object shapes and sizes, other models often fail due to reliance on object-specific features that may not generalize well to new instances. It highlights the potential for applications in dynamic, unstructured environments where robots encounter unfamiliar objects and must perform tasks with minimal human intervention.
4.5 Adapt to Real-World Bimanual Robot
Figure 8: (a) Environmental setup for the bimanual robot, featuring three camera views; (b) Table bussing setup, with a trash bin positioned on the right side and a panel on the left. The task requires placing all tableware on the panel and all trash in the trash bin; (c-f) All tableware and trash items used in the bussing table task evaluation.
In this section, we examine DiVLA’s adaptability to a dual-arm robot. Inspired by $\pi_{0}$ [7], we designed a table bussing task that involves clearing table with various objects. This task was adapted for a bimanual robot setup: all tableware should be placed on a panel to the left, while trash items are dropped into a bin on the right. Similar to our factory sorting task, we evaluated the model’s performance using both seen objects and a combination of seen and unseen objects. The environment setup, along with all objects used for training and evaluation, is shown in Figure 8. Our evaluation consisted of twelve trials, with three to five objects randomly placed on the table. The success rate is computed by how many objects are correctly placed.
Our experimental results show that our model successfully completes tasks in most cases when the object has appeared in the training data, achieving an average success rate of 72.9% on seen objects. In contrast, the Diffusion Policy and OpenVLA achieve 45.8% and 0% success rates. For tasks involving both seen and unseen objects, DiVLA achieves up to a 70.8% success rate, a slight decrease in success rate compared to the seen one, showing remarkable generalization to objects with different colors and shapes. Finally, DiVLA demonstrates the ability to recognize unseen objects, particularly by responding sensitively to object color. For example, it categorizes a Sprite can as a ”green can” and correctly places it in the trash bin. This observation further supports the idea that reasoning contributes to generalization.
Table 3: Experimental results for Table Bussing on real bimanual robot. Our method significantly outperforms both Diffusion Policy and OpenVLA by a large margin.
Scenarios Diffusion Policy OpenVLA DiVLA-2B
Seen 45.8 0 72.9
Mixed 31.2 0 70.8
4.6 Following Novel Instruction
The previous section primarily evaluated the model’s performance on in-distribution task commands. In this section, we assess the model’s ability to follow novel instructions, specifically focusing on its generalization to previously unseen commands. We introduce new instructions that prompt the model to
We tested the model on four objects: 1) a watermelon, 2) a lemonade, 3) a blue paper trash, and 4) a red pepper. This is an extremely challenging task, as these novel instructions are absent from both the Droid dataset and our collected data. We evaluated four new instructions, with results summarized in Table 2.
Our findings indicate that both OpenVLA and DiVLA-2B can recognize these unseen objects and perform basic pick-and-place tasks. However, when it comes to complex sequential tasks, OpenVLA fails to interpret the instructions accurately; instead, it randomly selects items. In contrast, our method correctly follows the instructions, picking up objects in the specified sequence. We hypothesize that by learning to decompose long tasks into subtasks, our method acquires a generalized ability to understand complex, multi-step instructions. While OpenVLA can execute simpler commands like ’Pick up the watermelon,’ it struggles with more advanced instructions requiring item selection in a specific order. Additionally, we observe a decrease in grasping precision when the model processes novel instructions, indicating that the novelty of instructions introduces further complexity to task execution.
5 Conclusion
In this work, we present DiVLA, a state-of-the-art vision-language-action model that delivers strong performance in both simulations and real-world scenarios, including single-arm and dual-arm robots. The core of our method lies in combining the next-token prediction objectives and diffusion models: the former for task reasoning and the latter for action prediction. We introduce a reasoning reuse module to enhance action generation and implement view-adaptive tokenization to reduce computational costs. Through extensive evaluations in both simulations and across multiple real-world embodiments, we demonstrate that DiVLA outperforms several SOTA robot models. Additionally, we show that DiVLA has robust generalization capabilities, adapting effectively to new instructions, tasks, and environments. Our research offers a novel perspective on designing VLA models, encouraging a rethinking of how reasoning can be reused to facilitate end-to-end policy learning.
Supplementary Material
Figure 9: Multi-task Learning and Visual Generalization. We evaluate each method on multi-task learning and visual generalizations, including adding additional distractors, changing the background, and implementing colorful lighting. Each set of three images represents the initial state, intermediate state, and final state of one trial.
Figure 10: Factory Sorting, Zero-Shot Bin-Picking and Table Bussing. We further evaluate all robot policies on additional challenging tasks, including factory sorting, zero-shot bin-picking, and bimanual table bussing. These tasks involve previously unseen objects with diverse textures, varying heights, and different degrees of deformability (as illustrated in Figure 6 and Figure 7). Each set of three images represents the progression of a single trial, showcasing the initial state, intermediate state, and final state.
Table 4: Summarization for the number of demonstrations and average trajectory length for our real-world tasks.
# Task # of Demonstrations Avg. Trajectory Length
1 Object selection 160 91
2 Flip the vertically placed pot 120 137
3 Place cube inside box with closed lid 50 146
4 Place cup onto the plate 100 107
5 Place cube into yellow/blue Box 100 90
6 Evaluation Tasks and Detailed Results
6.1 Evaluation Tasks
As described in Section 4, we evaluate our DiVLA and baselines on multi-task learning, factory sorting, zero-shot bin-picking, and table bussing. In this section, we provide details on these tasks. All of these tasks are depicted in Figure 9 and 10.
For multi-task learning, we evaluate all methods on 5 tasks for multi-task learning and 5 visual generalization tasks. We evaluate each method with a total of 77 trials for multi-task learning and 45 trials for visual generalization. We provide the number of demonstrations and the average length of these tasks in 4. A brief description of these tasks is as follows:
  • Object selection. In this task, we evaluate the model’s ability to comprehend and act on user intent effectively. To do so, we design three distinct instructions, each requiring the robot to identify and pick up the appropriate object based on the user’s expressed intent. These instructions are intentionally diverse to test the model’s versatility and robustness in understanding varying contexts and nuances of user commands.
  • Flip the vertically placed pot. In this task, we evaluate the model’s commonsense reasoning abilities. Specifically, the model must determine whether a vertically placed pot is oriented to the right or left and accurately identify its direction to position it correctly. This test assesses the model’s capacity to understand spatial orientation and apply logical reasoning to achieve the desired outcome.
  • Place cube inside box with closed lid. In this task, we evaluate the model’s ability to reason about and execute a multi-step operation that involves manipulating objects with physical constraints. The robot must first identify the box with the closed lid, open the lid, and then place the cube inside the box. This task assesses the model’s planning, understanding of spatial relationships, and ability to perform actions in a sequence to achieve a desired goal.
  • Place cup onto the plate. In this task, we test the model’s ability to perceive and reason about the spatial arrangement of objects in a structured environment. A plate is placed on a two-tiered shelf, and the robot must first identify which tier the plate is located on. And the robot needs to carefully place the cup onto the plate without disturbing the surrounding setup. This task evaluates the model’s spatial reasoning and precision in action.
  • Place cube into yellow/blue Box. In this task, we assess the model’s ability to follow instructions and demonstrate spatial reasoning. The robot is presented with an instruction specifying a designated box, either yellow or blue, and is required to accurately identify the correct box based on the given instruction. Once identified, the robot must then place a cube inside the chosen box. This task tests not only the model’s comprehension of language-based directives but also its ability to integrate spatial awareness and execute precise actions.
Setup for visual generalization. In this setting, we assess the model’s robustness and ability to generalize its visual perception to different environmental conditions. The robot is required to perform manipulation tasks under challenging visual variations, including randomly placed distractors, distracting lighting conditions, and a colorful background. These variations test the model’s capacity to maintain focus on the primary task while filtering out irrelevant visual noise and adapting to dynamic visual environments. The goal is to ensure the robot can accurately identify and interact with objects despite significant deviations from standard settings.
Moreover, detailed descriptions for factory sorting, zero-shot bin-picking, and table bussing are provided in Sections 4.3, 4.4, and 4.5, respectively.
Table 5: Detailed Experimental Results on both Franka and AgileX Aloha.
Type Task Trails DP TinyVLA Octo OpenVLA DiVLA-2B
Multi-Task Learning Object Selection 33 22 23 19 24 33
Flip the Vertically Placed Pot. 11 4 5 3 2 11
Place Cube into Yellow/Blue Box 11 0 3 3 6 10
Place Cup onto Plate 11 4 5 1 4 7
Place Cube inside Box with Closed Lip 11 0 4 2 2 7
Visual Generalization Object Selection 9 1 4 4 5 4
Flip the Vertically Placed Pot. 9 1 4 1 1 6
Place Cube into Yellow/Blue Box 9 0 2 0 3 6
Place Cup onto Plate 9 2 2 0 3 6
Place Cube inside Box with Closed Lip 9 0 1 2 0 4
More Challenging Tasks Factory Sorting (Seen) 80 27 39 33 45 61
Factory Sorting (Mixed) 80 11 25 27 36 50
Factory Sorting (Cluttered Seen) 65 16 21 17 30 43
Factory Sorting (Cluttered Mixed) 65 6 14 15 22 40
Zero-Shot Bin-Picking 102 10 24 4 29 65
Table Bussing (Seen) 48 22 - - 0 35
Table Bussing (Mixed) 48 15 - - 0 34
Table 6: Inference speed for DiVLA. We report the inference speed for DiVLA on A6000 GPU.
Method DiVLA-2B DiVLA-7B OpenVLA-7B
Control Frequency 82Hz 42Hz 5Hz
Figure 11: View Generalization. We evaluate DP, OpenVLA, and DiVLA-2B in the view shifting setting, where we use completely different camera positions to capture images. The blue part indicates the original camera positions, while the red part indicates the new camera positions.
6.2 Detailed Results
In this section, we present the complete evaluation results for both the Franka and Bimanual AgileX robots, as detailed in Table 5. DiVLA-2B consistently demonstrates superior performance across the majority of tasks. Additionally, our findings highlight DiVLA-2B’s robust visual generalization capabilities, effectively adapting to changes in the surrounding environment.
Table 7: Experimental results for view shifting generalization. We conduct view shifting generalization on factory sorting task. The experimental setup is shown in Figure 11. We report the success rate for each policy. Each method is evaluate with 5 trails.
Task $\setminus$ Method Diffusion Policy OpenVLA DiVLA-2B
Sorting 0 0 60
7 Implementation Details
For the baselines, we follow a uniform training strategy to ensure consistency. For OpenVLA, we follow the original implementation which only uses a single camera. We also test OpenVLA with three camera views, where each view goes through the same visual encoder and concatenates its outputs for further process. We observe some performance improvement over multiple camera views. For instance, for sorting task, the average success rate will increase from 32.3% to 45.3%. Nevertheless, since most of the existing works only report the OpenVLA baseline with a single camera view, we only report experimental results for this setting. More details can refer to TinyVLA [57]. OpenVLA’s pre-trained weights are used, and we fine-tune the model on our dataset with a learning rate of 2e-5. For the Diffusion Policy, we employ DistilBERT to encode language instructions, adopting a methodology similar to YAY [44]. All models are trained for the same number of epochs, and only the final checkpoint is used for evaluation. To ensure a fair comparison, we avoid cherry-picking the checkpoint.
8 More Experiments
8.1 View Shifting Generalization
Imitation learning often struggles to generalize its capabilities to novel viewpoints. In this study, we evaluate the view generalization performance of DP, OpenVLA, and DiVLA-2B in a factory sorting task. The experimental setup is illustrated in Figure 11. As shown in Table 7, OpenVLA exhibits poor generalization to view shifts. In contrast, DiVLA-2B demonstrates notable robustness in view generalization, achieving a success rate of 60%. This result highlights the advantages of leveraging pretraining on large-scale robotic datasets to enhance generalization capabilities.
8.2 Efficient Inference
Fast inference speed is critical for deploying VLA models in real-world applications. However, as models grow in size and complexity, their inference speed tends to decrease significantly on server-grade hardware. To address this challenge and enhance model performance, we deploy our models on the vLLM framework [27]. This approach leverages optimized infrastructure to maximize inference efficiency. In Table 6, we present a comparative analysis of control frequencies for DiVLA-2B, DiVLA-7B, and OpenVLA (a 7B model). Our results demonstrate remarkable speed improvements: the DiVLA-2B model achieves an impressive 82 Hz on an A6000 GPU, showcasing exceptional performance. Similarly, DiVLA-7B achieves a control frequency of 42 Hz, which is 8 times faster than OpenVLA at the same model size. These findings underscore the effectiveness of our optimizations in maintaining scalability without sacrificing speed, paving the way for broader real-world applicability of VLA models.
Notably, we avoid using action chunking techniques employed in other works, such as Data-Scaling-Laws [30] and $\pi_{0}$ [7], as we find that while action chunking can improve inference speed, it adversely impacts generalization performance. Another key observation is that our model experiences significant performance degradation when evaluated at lower bit precisions, such as 8-bit and 4-bit. This aligns with findings in OpenVLA, which highlight performance fluctuations introduced by quantization. These results suggest that current VLA models may require specifically designed quantization methods to maintain performance while achieving fast inference speeds at low precision.
8.3 Model Scaling
An important consideration for pure end-to-end neural networks is assessing their ability to scale effectively. As the model size and data volume grow, the performance of the model should ideally improve. To evaluate this scalability, we conducted additional experiments on two challenging tasks: factory sorting and zero-shot bin-picking. These tasks provide rigorous benchmarks to measure how increased model capacity and pre-training data influence performance. The detailed experimental results are presented in Table 8.
Our analysis highlights significant performance improvements when larger models and more extensive pre-training datasets are utilized. For example, the DiVLA-7B model demonstrates substantial gains over the DiVLA-2B model, achieving improvements of 8.7% and 3.0% in success rates for the factory sorting and bin-picking tasks, respectively. This underscores the benefits of scaling model size, which enables better representation learning and decision-making capabilities.
Moreover, scaling up to 72B parameters yields even greater performance boosts. These larger models not only achieve higher success rates in in-distribution scenarios but also exhibit superior generalization to out-of-distribution setups. This improvement reflects the models’ enhanced ability to handle variability in task environments, unseen object properties, and dynamic conditions.
Table 8: Experimental results for model scaling.
Tasks/Models DiVLA-2B DiVLA-7B DiVLA-72B
Sorting 66.2 74.9 82.4
Bin Picking 63.7 66.7 75.9
Table 9: VQA for DiVLA. We test DiVLA’s ability to answer questions based on visual signals.
Question Object/Scene Answer Correct
What is the object? [Uncaptioned image] Toy Tiger
[Uncaptioned image] Tulip
[Uncaptioned image] Orange
What is the object’s color? [Uncaptioned image] Brown
[Uncaptioned image] Blue
[Uncaptioned image] Green
Describe the scene. [Uncaptioned image] The cube is on the right side of yellow pepper
[Uncaptioned image] The ball is on the top of a holder
8.4 DiVLA can do Visual-Question-Answering
Previous works, such as RT-2 [9] and ECoT [64], have suggested that co-training with vision-language data helps preserve basic conversational functionality for Vision-Language Action (VLA) data. In this section, we demonstrate that despite not being co-trained with vision-language data, our method retains its chatting capability and is proficient in performing common visual-question-answering (VQA) tasks. However, these works have neither provided concrete examples of chat interactions nor open-sourced their results, leaving it unclear to what extent conversational capabilities are retained in such models. We give a few examples in Table 9.
When prompted to identify objects, we observe that the model demonstrates the ability to recognize some items accurately, such as tulips and oranges. However, it struggles with more nuanced distinctions; for instance, DiVLA mistakenly identifies a toy dragon as a toy tiger, likely due to its reliance on color features rather than other distinguishing characteristics like shape or texture.
Interestingly, we observe that the model demonstrates strong sensitivity to color, consistently providing correct answers to all three questions focusing on an object’s color. This highlights a notable strength in color recognition but also reveals limitations in processing broader visual features for complex tasks. To further assess its capabilities, we tested the model with a simple scene description task, requiring it to describe spatial relationships between objects. The model successfully interpreted spatial relationships, such as identifying that one object is on the right or on top of another. However, it often failed to recognize objects correctly. For example, DiVLA misclassified a football as a regular ball.
Notably, these objects were not included in the model’s pre-training or fine-tuning data. However, the pre-trained VLM may have encountered similar objects during its pre-training phase, enabling it to recognize them correctly in some cases. This observation underscores the importance of leveraging pre-trained VLMs as a foundation for end-to-end visuomotor learning, as they provide a strong prior understanding of visual concepts that can significantly enhance downstream performance in complex tasks.